OpenStudy (anonymous):

If 5^-2 = 1/5^2, why does the neg exponent change to a positive??? (Its been a very long time since ive been in any math class)

OpenStudy (anonymous):

Because you put it on the bottom of a fraction basically its just the definition.

OpenStudy (anonymous):

k. thx

OpenStudy (mathstudent55):

Think of it this way: 32 = 2^5 = 2 x 2 x 2 x 2 x 2 16 = 2^ 4 = 2 x 2 x 2 x 2 8 = 2^3 = 2 x 2 x 2 4 = 2^2 = 2 x 2 2 = 2^1 = 2 1 = 2^0 = 1 Look at the number on left, 32, 16, 8, 4, ... To go from one line to the next you divide by two. The middle number, 2^5, 2^4, has an exponent 1 less than the line before. So do it again for the next few steps: 1/2 = 2^(-1) = 1/2 1/4 = 2^(-2) = 1/(2 x 2) etc.

OpenStudy (mathstudent55):

It's what DarkGhost said, just the definition, but the definition of negative exponents was created on purpose to make it work.